OpenStudy (anonymous):

A cone has its tip at the point (0,0,5) and its base the disk D, x^2+y^2<=4, in the xy-plane. The surface of the cone is the curved and slanted face, S, oriented upward, and the flat base, D, oriented downward. The flux of the constant vector field F=ai+bj+czk through S is given by integral(of S) of F dA =5.48. What is the integral(of D) of F dA equal to? I'm stuck :(

OpenStudy (anonymous):

Lol, any help on this is much appreciated. Calc III hw is killing me

OpenStudy (anonymous):

hmm, I don't remmember exactly, but i thin that the flux of constant field over closed volume shouldn't be 0? That would give -5.48 for the D part

OpenStudy (anonymous):

If F = <a, b, c> then the flux over just D would be -5.48. That was the first part of the question. The second part states that F now equals <a, b, cz> so it's not techincally constant, and I apologize for the wording.

OpenStudy (anonymous):

ok... np. let me think, :)

OpenStudy (anonymous):

The Field is not constant

OpenStudy (anonymous):

z is perpendicular to D. So just simple integral in 1 variable

OpenStudy (anonymous):

Is there a misprint?

OpenStudy (anonymous):

myko....what would that integral be then in this case?

OpenStudy (anonymous):

Is the field F=ai+bj+czk or F=ai+bj+ck

OpenStudy (anonymous):

eliasaab...F=<a, b, cz> as was stated in my reply a few moments ago. Sorry for the confusion.

OpenStudy (anonymous):

Great, that makes it easy. Stay tuned.

OpenStudy (anonymous):

What is the volume of your cone? DO you know hoe to do that?

OpenStudy (anonymous):

Well the the tip of the cone is (0,0,8) and the base is in the form of x^2 + y^2 <=9, so V=(1/3)(pi)*(r^2)*h where r=3 and h=8?

OpenStudy (anonymous):

\[ V=\frac{1}{3} \pi h r^2 \]

OpenStudy (anonymous):

r =2 and h=5

OpenStudy (anonymous):

Err, yeah. The numbers on the problem I posted are different from my actual problem on the homework. So yeah, lets just go with r=2, h=5

OpenStudy (anonymous):

What's the divergence of your field?

OpenStudy (anonymous):

divF = c am i right?

OpenStudy (anonymous):

\[fluxD = \int\limits \int\limits _{S}zkdxdy\] to solve this parametrize the circle

OpenStudy (anonymous):

myko, where did u get the integrand "zk dx dy" ?

OpenStudy (anonymous):

Yes Flux of your field \[ Flux= \int \int \int_{Cone} \text { Divergence (F) } dV \]

OpenStudy (anonymous):

I did try doing what you just did, eliassab, but that's just giving my the flux from the cone, which is in this case 5.48. I'm guessing solve for c from that integral?

OpenStudy (anonymous):

That is what i got, yes..

OpenStudy (anonymous):

Since you know c, it easy to compute the flux around the base.

OpenStudy (anonymous):

Your filed is now known. Apply the definition of the flux and you are done

OpenStudy (anonymous):

across the base

OpenStudy (anonymous):

Flux at the base = c* pi(r^2)?

OpenStudy (anonymous):

- c pi r^2 since n=(0,0,-1}

OpenStudy (anonymous):

Oh okay, let me put this in to my web hw and see if it's all good.

OpenStudy (anonymous):

Input the right numbers and not the numbers posted.

OpenStudy (anonymous):

so I did try doing this, where in my case h = 8 and r = 3

OpenStudy (anonymous):

solving for c, I got 3.35, and thus the flux at the base should've been -c*pi*3^2

OpenStudy (anonymous):

giving me -94.719

OpenStudy (anonymous):

but the answer is wrong...

OpenStudy (anonymous):

Repost the the exact problem and let me see.

OpenStudy (anonymous):

A cone has its tip at the point (0,0,8) and its base the disk D, x^2+y^2<=9, in the xy-plane. The surface of the cone is the curved and slanted face, S, oriented upward, and the flat base, D, oriented downward. The flux of the t vector field F=ai+bj+czk through S is given by integral(of S) of F dA =3.12. What is the integral(of D) of F dA equal to?

OpenStudy (anonymous):

I got 3.3518

OpenStudy (anonymous):

then according to the equation you gave me where (1/3)c(pi)(h) = r^2(flux given)

OpenStudy (anonymous):

Lol, so c = .01379 about?

OpenStudy (anonymous):

so c= .04138 ?

OpenStudy (anonymous):

then -c(pi*r^2) gives the flux through D, where c=.04138?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, let me try

OpenStudy (anonymous):

it is about -1.17

OpenStudy (anonymous):

Hmm, still saying it's wrong.

OpenStudy (anonymous):

I know what went wrong. Stay tuned.

OpenStudy (anonymous):

The 3.12 is the flux thru S and not thru the entire outside area of the cone. Hence \[ 3.12 -\pi c r^2=c \frac{1}{3} \pi h r^2 \] Solve for c, you get \[ c\approx .03 \] your answer is \[ -c \pi r^2 =-0.85 \]